OpenStudy (anonymous):

-5√192) - 2 √243) - 8√108) subtract terms

OpenStudy (anonymous):

\[-5\sqrt(64*3)-2\sqrt(81*3)-8\sqrt(36*3)\] \[=(-5*8*\sqrt3)-(2*9*\sqrt3)-(8*6*\sqrt3)\] \[=\sqrt3(-40-18-48)\] \[=-106\sqrt3\]